FreeStatisticalTools.com
Statistical tables

Z Table: Standard Normal Distribution Table

Cumulative probabilities for positive and negative z scores, critical z values for confidence intervals, and a calculator that converts between z scores and probabilities.

z from −3.49 to 3.494 decimal placesPrintable
Standard normal cumulative probabilities for positive z
z0.000.010.020.030.040.050.060.070.080.09
0.00.50000.50400.50800.51200.51600.51990.52390.52790.53190.5359
0.10.53980.54380.54780.55170.55570.55960.56360.56750.57140.5753
0.20.57930.58320.58710.59100.59480.59870.60260.60640.61030.6141
0.30.61790.62170.62550.62930.63310.63680.64060.64430.64800.6517
0.40.65540.65910.66280.66640.67000.67360.67720.68080.68440.6879
0.50.69150.69500.69850.70190.70540.70880.71230.71570.71900.7224
0.60.72570.72910.73240.73570.73890.74220.74540.74860.75170.7549
0.70.75800.76110.76420.76730.77040.77340.77640.77940.78230.7852
0.80.78810.79100.79390.79670.79950.80230.80510.80780.81060.8133
0.90.81590.81860.82120.82380.82640.82890.83150.83400.83650.8389
1.00.84130.84380.84610.84850.85080.85310.85540.85770.85990.8621
1.10.86430.86650.86860.87080.87290.87490.87700.87900.88100.8830
1.20.88490.88690.88880.89070.89250.89440.89620.89800.89970.9015
1.30.90320.90490.90660.90820.90990.91150.91310.91470.91620.9177
1.40.91920.92070.92220.92360.92510.92650.92790.92920.93060.9319
1.50.93320.93450.93570.93700.93820.93940.94060.94180.94290.9441
1.60.94520.94630.94740.94840.94950.95050.95150.95250.95350.9545
1.70.95540.95640.95730.95820.95910.95990.96080.96160.96250.9633
1.80.96410.96490.96560.96640.96710.96780.96860.96930.96990.9706
1.90.97130.97190.97260.97320.97380.97440.97500.97560.97610.9767
2.00.97720.97780.97830.97880.97930.97980.98030.98080.98120.9817
2.10.98210.98260.98300.98340.98380.98420.98460.98500.98540.9857
2.20.98610.98640.98680.98710.98750.98780.98810.98840.98870.9890
2.30.98930.98960.98980.99010.99040.99060.99090.99110.99130.9916
2.40.99180.99200.99220.99250.99270.99290.99310.99320.99340.9936
2.50.99380.99400.99410.99430.99450.99460.99480.99490.99510.9952
2.60.99530.99550.99560.99570.99590.99600.99610.99620.99630.9964
2.70.99650.99660.99670.99680.99690.99700.99710.99720.99730.9974
2.80.99740.99750.99760.99770.99770.99780.99790.99790.99800.9981
2.90.99810.99820.99820.99830.99840.99840.99850.99850.99860.9986
3.00.99870.99870.99870.99880.99880.99890.99890.99890.99900.9990
3.10.99900.99910.99910.99910.99920.99920.99920.99920.99930.9993
3.20.99930.99930.99940.99940.99940.99940.99940.99950.99950.9995
3.30.99950.99950.99950.99960.99960.99960.99960.99960.99960.9997
3.40.99970.99970.99970.99970.99970.99970.99970.99970.99970.9998

Each cell is the area under the standard normal curve to the left of z, P(Z < z). Row = z to one decimal; column = second decimal.

Values computed with SciPy and checked against published tables. Free to use and print for teaching. Please link to this page if you share it.

Critical z values for confidence intervals and tests

Confidence levelTwo-tailed zOne-tailed z
80%1.2820.842
90%1.6451.282
95%1.9601.645
98%2.3262.054
99%2.5762.326
99.9%3.2913.090

How to read the z table

  1. Split your z score into the first decimal (row) and the second decimal (column). For z = 1.96, the row is 1.9 and the column is 0.06.
  2. The cell where they meet, 0.9750, is the proportion of the distribution below z = 1.96.
  3. For the area above z, subtract from 1: 1 − 0.9750 = 0.0250. For a two-tailed p-value, double the smaller tail: 2 × 0.0250 = .05.

Worked example

Exam scores have a mean of 70 and SD of 8. What share of students scored below 82? First convert to a z score: z = (82 − 70) / 8 = 1.50. The table gives 0.9332 for z = 1.50, so about 93.3% of students scored below 82 and 6.7% scored above.

Negative z scores

Use the negative table, or use symmetry: P(Z < −z) = 1 − P(Z < z). For example P(Z < −1.50) = 1 − 0.9332 = 0.0668.

z = (x − μ) / σ

Frequently asked questions

What is the z value for 95% confidence?

1.96 for a two-sided interval or two-tailed test. The one-tailed value at α = .05 is 1.645.

What does the value in a z table mean?

It is the cumulative probability: the area under the standard normal curve to the left of the z score, which is the proportion of values below that point.

How do I find the area between two z scores?

Look up both values and subtract the smaller cumulative probability from the larger. For example, between z = −1 and z = 1 the area is 0.8413 − 0.1587 = 0.6827, about 68%.

What is the probability for z = 2?

P(Z < 2.00) = 0.9772, so 2.28% of values lie above z = 2 and about 95.45% lie between −2 and 2.

When should I use a t table instead of a z table?

Use the t table when the population standard deviation is unknown and estimated from the sample, which is almost always the case with real data and small samples. The z table applies to standardized scores and large-sample tests of proportions.

Related tools